OpenStudy (anonymous):

Brain weight B as a function of body weight W in fish has been modeled by the power function B=.007W2/3, where B and W are measured in grams. A model for body weight as a function of body length L (measured in cm) is W=.12L2.53. If, over 10 million years, the average length of a certain species of fish evolved from 15cm to 20cm at a constant rate, how fast was the species' brain growing when the average length was 18cm? Round your answer to the nearest hundredth.

OpenStudy (anonymous):

where r u stuck at? do you have an equation made up?

OpenStudy (anonymous):

like i don't even know what equation to use, that's the problem

OpenStudy (anonymous):

oh haha I get that way too, sooo....

OpenStudy (anonymous):

so here are all that we know: B=.007W2/3 W=.12L2.53 15cm to 20cm 18cm [find]

OpenStudy (anonymous):

do you see how it says, "constant rate" it means slope of a line, what class is this? algebra?

OpenStudy (anonymous):

are you guys going over the 'line equation'?

OpenStudy (anonymous):

B=.007(.12L2.53)2/3 y=.007(.12x2.53)2/3 m = (y-y)/(x-x) find the slope, plug in the 15 and 20 in for x and get y, then find the slop[e